OpenStudy (anonymous):

Simplify (4/9)^-1/2 A. 2/9 B. 2/3 C. 4/3 D. 3/2

OpenStudy (anonymous):

a

OpenStudy (kropot72):

\[\large (\frac{4}{9})^{-\frac{1}{2}}=\frac{1}{\sqrt{\frac{4}{9}}}\]

OpenStudy (anonymous):

So b then?

Directrix (directrix):

Not b. @AlexisSM Try again.